# Pattern and Origin for the Extreme -ray Flares of 3C 454.3 and 3C 279: An Astrophysical Critical Damper?

### v1

@@ 5 Summary
Based on our understanding of the non-thermal radiation variability characteristics of blazars (including the long-term stochastic variability and QPOs: Zhang etĀ al., 2021, 2022, 2023), we further investigate the properties of the extreme -ray flares of 3C 454.3 and 3C 279. The GP method is carried out with SHO, DRW and Matrn as the kernels for the flaring data. The short-term variability modes for both 3C 454.3 flare and 3C 279 flare are more inclined towards the over-damped SHO mode (approaching a critically damped mode) and Matrn model. In particular, the systems producing the giant flares could be described by a critical damper. We build a general physical scenario to interpret our results and propose that the process in the jet that triggers abrupt energy release (e.g., magnetic reconnection) could be the cause of the two giant flares. The derived timescale of the variability provides a constraint on the size of the magnetic reconnection region.
@@ acknowledgments
D.H.Y acknowledges funding support from the National Natural Science Foundation of China (NSFC) under grant No. 12393852. We thank the support from the Postdoctoral Fellowship Program of China Postdoctoral Science Foundation (CPSF) under Grant Number GZB20230618.

### v2

@@ 6 Summary
Based on our understanding of the non-thermal radiation variability characteristics of blazars (including the long-term stochastic variability and QPOs; Zhang etĀ al., 2021, 2022, 2023), we further investigate the properties of the extreme -ray flares of 3C 454.3 and 3C 279. The GP method is carried out with SHO, DRW and Matrn as the kernels for the flaring data. The short-term variability modes for both 3C 454.3 flare and 3C 279 flare are more inclined towards the over-damped SHO mode (approaching a critically damped mode) and Matrn model. We propose that the variability of the giant -ray flares may reflect physical processes (in jets) analogous to a critically damped response. In such a process/system, magnetic reconnection may be the energy dissipative mechanism, resulting abrupt energy release. The derived timescale of the variability may provide an indication of the size of the magnetic reconnection region.
@@ acknowledgments
We acknowledge the funding support from the National Natural Science Foundation of China (NSFC) under grant No. 12393852, and the support from the Postdoctoral Fellowship Program of China Postdoctoral Science Foundation (CPSF) under Grant Number GZB20230618.
Appendix A Additional Figures
We examined the impact of parameter autocorrelation and different sampling methods on inference results. As an example, FigureĀ 5 shows the posterior distributions of the SHO model parameters fitted to the flare of 3C 454.3, obtained using both thinned MCMC samples and the nested sampling method.